Perimeter of a rectangle. The width of a rectangular playground is 2x – 5 ft., and the length is 3x + 9 ft. Write a polynomial P(x) that represents the perimeter and then evaluate this perimeter polynomial if x is 4 ft.
Answer by bcpm36(2)

P = 2(l + w) l = 3x + 9 ft w = 2x - 5 ft
P = 2( 3x + 9 + 2x - 5 )
p = 2 (5x + 4)
p = 10x + 8
x = 4
P = 10(4) + 8
P = 40 + 8
P = 48 ft